solve this system of equations by graphing.
y=3x-8
y=x-8

Answer by richwmiller(17219)

y=3x-8
y=x-8
solve for intercepts where x=0 and y=0
0=3x-8
8=3x
8/3=x
2.67=x
y=3(0)-8
y=-8
y=x-8
0=x-8
8=x
y=0-8=-8
plot those four points and draw the appropriate lines